Find an equation of the line that contains the given point and is parallel to the given line: 2x-y=5, (-3,-8)
Found 2 solutions by richwmiller, stanbon:
Answer by richwmiller(17219)

parallel means same slope
slope is -a/b=-2/-1=2
or 2x-y=5
2x-5=y
m=2
y=mx+b
-8=2*(-3)+b
-8=-6+b
-2=b
y=2x-2

Answer by stanbon(75887)

Find an equation of the line that contains the given point and is parallel to the given line: 2x-y=5, (-3,-8)
--------------------------------
Slope-intercept:
y = 2x-5
slope = 2
-----------
Every line that is parallel must have slope = 2.
----------------------
Form of the equation you want: y = mx+b
You know m = 2 and y = -8 when x = -3
-----
Solve for "b":
-8 = 2(-3)+b
-8 = -6 + b
b = -2
---------------------
Equation:
y = 2x-2
======================
Cheers,
Stan H.